Use the following functions that show quantity demanded and quantity supplied to answer the question

Qd=600 - 30P

Qs=-300 +120P

If the price is R8, then there will be a surplus of ______ units.


1
Expert's answer
2021-06-01T12:12:41-0400

Finding the surplus we have to replace the price into both the supply and demand functions and find the difference between the two.

"Qd=600-30(8)=360"

"Qs= -300+120(8)=660"

"Qs-Qd = 660-360"

"Surplus = 300"